# Shear flow of frictional spheroids: lentil-like grains slide more than rice-like ones

### v1

@@ Appendix D Normal stress difference
Looking at Figure 15 the first normal stress difference shows an increasing trend with $I$ for $\mu_{p}=0$ and a slightly decreasing one in the frictional case. We report $N_{1}<0$ at $I>10^{-2}$ for slightly oblate particles $0.830.6<\alpha<$ 0.83 at $\mu_{p}=0.4$, even though no shear band is observed.
@@
Figure 15: First (top) and second (bottom) normalized normal stress differences as a function of inertial number. Left $\mu_{p}=0$, middle $\mu_{p}=0.4$, right $\mu_{p}=10.0$. Same color legend as in Figure 6.
@@ Appendix E Particle circulation and diffusive behaviour
The diffusion coefficient is extracted from the mean squared displacement using the Einstein relation, $D_{k}=\mathrm{lim}_{t\to\infty}\langle k^{2}\rangle/2t$, with $k=y,z$, and numerically estimated via least-square fitting. As shown in Figure 16, for frictionless particles, the diffusion coefficient exhibits a slightly decreasing trend with the inertial number, while for frictional cases, it remains nearly constant. Unlike previous studies [53, 54, 61], our data do not follow the expected scaling $D_{yy}\sim\dot{\gamma}d^{2}I^{-1/2}$, likely due to finite size effects, where approximately 15 particles span the domain in the $y-$direction, constraining vortex formation to the system size. Even in the case of frictionless spheres, the best-fit power law decay is around $D_{yy}\sim\dot{\gamma}d^{2}I^{-1/8}$. Another intriguing avenue to explore is whether the observed reduction is a distinctive feature of 3D systems—analogous to vortex stretching in Newtonian fluids, which occurs exclusively in three dimensions.
Despite some noise in the data, elongated and flat particles generally exhibit lower diffusion compared to spheres in all cases except for the highly collisional-frictional regime. In most cases, the shear-induced alignment of elongated particles creates a cage-like structure that restricts their mobility, thereby reducing diffusion [18]. Interestingly, oblate particles exhibit greater diffusion in the vorticity direction ($z$) compared to prolate ones, suggesting a weaker caging effect along this axis.
For frictional particles, we also observe entanglement and rotational motion, similar to what was reported for flat disks and elongated rods under Lees-Edwards boundary conditions [16]. In our simulations, where simple shear is achieved by physically tilting the simulation box, this phenomenon primarily occurs when the box tilt is near zero (see movie in supplemental material. The circulation arises due to the absence of box tilt, allowing particles to move coherently along $y$ since they are reinserted at approximately the same $x$-coordinate when crossing the periodic boundary. This effect is further supported by the mean square displacement in $y$, which exhibits periodic peaks over time (data not shown). Additional simulations with $L_{x}>7L_{y}$ confirm a similar behavior. Furthermore, the observed coherent motion is linked to the vortex structures described in [54], warranting further investigation.
Figure 16: Gradient (top) and vorticity (bottom) diffusion coefficient, normalized by $\dot{\gamma}d_{eq}^{2}$, as a function of inertial number. Left $\mu_{p}=0$, middle $\mu_{p}=0.4$, right $\mu_{p}=10.0$. Same color legend as in Figure 6.
@@ Appendix F Spatial distribution of contacts
Figure 17: Angular distribution of contacts normal (a) direction in the global reference frame, i.e. angle$=0^{\circ}$ for a contact along $x$, $90^{\circ}$ for a contact along $y$. Normalized conditional force $\overline{F}_{n}/PA_{p}$ (b) and $\overline{F}_{t}/PA_{p}$ (c). Same color legend as in Figure 6.
@@ Appendix G Decomposition of shear stress in normal and tangential component
The tangential component of the shear stress generally increases as $\mu_{p}$ is increased, and becomes dominant at $|r_{g}|>0.33$ higher values for both flat and elongated particles. The total shear stress and its tangential component are maximum for flat particles at $\mu_{p}\approx 1.0$, and not in the high friction limit. We therefore conclude that, given a certain shape, the maximum shear strength is found at finite values of $\mu_{p}$, those for which the alignment of the optimal particles is achieved.

### v2

@@ Appendix D Normal stress difference
Looking at Figure 15 the first normal stress difference shows an increasing trend with $I$ for $\mu_{p}=0$ and a slightly decreasing one in the frictional case. We report $N_{1}<0$ at $I>10^{-2}$ for slightly oblate particles $0.6<\alpha<0.83$ at $\mu_{p}=0.4$, even though no shear band is observed.
@@
Figure 15: First (top) and second (bottom) normalized normal stress differences as a function of inertial number.
@@ Appendix E Particle circulation and diffusive behaviour
The diffusion coefficient is extracted from the mean squared displacement using the Einstein relation, $D_{k}=\mathrm{lim}_{t\to\infty}\langle k^{2}\rangle/2t$, with $k=y,z$, and numerically estimated via least-square fitting. As shown in Figure 16, for frictionless particles, the diffusion coefficient exhibits a slightly decreasing trend with the inertial number, while for frictional cases, it remains nearly constant. Unlike previous studies [61, 62, 72], our data do not follow the expected scaling $D_{yy}\sim\dot{\gamma}d^{2}I^{-1/2}$, likely due to finite size effects, where approximately 15 particles span the domain in the $y$-direction, constraining vortex formation to the system size. Even in the case of frictionless spheres, the best-fit power law decay is around $D_{yy}\sim\dot{\gamma}d^{2}I^{-1/8}$. It is likely that the observed reduction is a distinctive feature of 3D systems—analogous to vortex stretching in Newtonian fluids, which occurs exclusively in three dimensions [69].
Despite some noise in the data, elongated and flat particles generally exhibit lower diffusion compared to spheres in all cases except for the highly frictional rolling regime. In most cases, the shear-induced alignment of elongated particles creates a cage-like structure that restricts their mobility, thereby reducing diffusion [18]. Interestingly, oblate particles exhibit greater diffusion in the vorticity direction ($z$) compared to prolate ones, suggesting a weaker caging effect along this axis.
For frictional particles, we also observe entanglement and rotational motion, similar to what was reported for flat disks and elongated rods under Lees-Edwards boundary conditions [16]. In our simulations, where simple shear is achieved by physically tilting the simulation box, this phenomenon primarily occurs when the box tilt is near zero (see movie in supplemental material [64]). The circulation arises due to the absence of box tilt, allowing particles to move coherently along $y$ since they are reinserted at approximately the same $x$-coordinate when crossing the periodic boundary. This effect is further supported by the mean square displacement in $y$, which exhibits periodic peaks over time (data not shown). Additional simulations with $L_{x}>7L_{y}$ confirm a similar behavior. The observed coherent motion is linked to the vortex structures described in [62], which in 3D can span the whole domain size at high densities, unlike the correlation lengths [69].
Figure 16: Gradient (top) and vorticity (bottom) diffusion coefficient, normalized by $\dot{\gamma}d_{eq}^{2}$, as a function of inertial number.
@@ Appendix G Decomposition of shear stress in normal and tangential component
The tangential component of the shear stress generally increases as $\mu_{p}$ is increased, and becomes dominant at $|r_{g}|>0.33$ higher values for both flat and elongated particles. The total shear stress and its tangential component are maximum for flat particles at $\mu_{p}\approx 1.0$, and not in the high friction limit. We therefore conclude that, given a certain shape, the maximum shear strength is found at finite values of $\mu_{p}$, those for which the optimal alignment of particles is achieved.